# InterControl: Zero-shot Human Interaction Generation by Controlling Every Joint

### v3

@@ Abstract
Text-conditioned motion synthesis has made remarkable progress with the emergence of diffusion models. However, the majority of these motion diffusion models are primarily designed for a single character and overlook multi-human interactions. In our approach, we strive to explore this problem by synthesizing human motion with interactions for a group of characters of any size in a zero-shot manner. The key aspect of our approach is the adaptation of human-wise interactions as pairs of human joints that can be either in contact or separated by a desired distance. In contrast to existing methods that necessitate training motion generation models on multi-human motion datasets with a fixed number of characters, our approach inherently possesses the flexibility to model human interactions involving an arbitrary number of individuals, thereby transcending the limitations imposed by the training data. We introduce a novel controllable motion generation method, InterControl, to encourage the synthesized motions maintaining the desired distance between joint pairs. It consists of a motion controller and an inverse kinematics guidance module that realistically and accurately aligns the joints of synthesized characters to the desired location. Furthermore, we demonstrate that the distance between joint pairs for human-wise interactions can be generated using an off-the-shelf Large Language Model (LLM). Experimental results highlight the capability of our framework to generate interactions with multiple human characters and its potential to work with off-the-shelf physics-based character simulators.
@@ 2.3 Controllable Diffusion Models
Diffusion-based generative models have achieved great progress in generating various modalities, such as image [50, 22, 9, 53], video [11, 17, 24] and audio [32]. Conditions and controlling ability in diffusion models are also well studied: (1) Inpainting-based methods [8, 7] predict part of the data with the observed parts as condition and rely on diffusion model to generate consistent output, which is used in PriorMDM [51]. (2) Classifier-guidance [9] trains a separate classifier and exploits the gradient of classifier to guide the diffusion process. Our InterControl inherits the spirit of classifier-guidance, yet our guidance is provided by Inverse Kinematics (IK) and no classifier is needed. (3) Classifier-free guidance [22] trains a conditional and an unconditional diffusion model simultaneously and trade-off its quality and diversity by setting weights. (4) ControlNet [70] introduces a trainable copy of pretrained diffusion model to process the condition and freezes the original model to avoid degeneration of generation ability. It enables diverse types of dense control signals for various purpose with minimal finetuning effort. Our InterControl also incorporate the idea of ControlNet [70] to finetune the pretrained MDM [55] to process spatial control signals and imporve the quality of generated motions after joint control.
@@ 3.1 Formulation of Interaction Generation
Inspired by human-scene interaction [64], we define human interactions as joint contact pairs $\mathcal{C}=\left\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots\right\}$, where $\mathcal{S}_{i}$ is the $i^{th}$ contact step. Taking two-person interaction as an example, each step $\mathcal{S}$ has several contact pairs $\mathcal{S}=\left\{\left\{j^{1}_{1},j^{2}_{1},t^{s}_{1},t^{e}_{1},c_{1},d_{1}\right\},\left\{j^{1}_{2},j^{2}_{2},t^{s}_{2},t^{e}_{2},c_{2},d_{2}\right\},\ldots\right\}$, where $j^{1}_{k}$ is the joint of person 1, $j^{2}_{k}$ is the joint of person 2, $t^{s}_{k}$ and $t^{e}_{k}$ means the start and end frame of the interaction, $c_{k}$ means contact type from {contact, avoid} to pull or push the joint pairs, $d_{k}$ is the desired distance in the interaction. By converting the contact pairs $\mathcal{S}$ to the mask $\boldsymbol{m}$ and distance $d$, and taking others’ joint positions as condition, we could guide the multi-person motion generation process to interact between joints in the form of spatial distance. In this way, interaction generation is transformed to be controllable single-person motion generation taking a text prompt $\boldsymbol{p}$ and a spatial control signal $\boldsymbol{c}\in\mathbb{R}^{N\times J\times 3}$ as input. Its goal is to predict motion sequence $\boldsymbol{x}\in\mathbb{R}^{N\times D}$ whose joints in the global space is aligned with spatial control $\boldsymbol{c}$, where $N$ is number of frames, $J$ is number of joints (e.g., 24 in SMPL [38]), and $D$ is the dimension of relative joint representations (e.g., 263 in HumanML3D [14]). It is non-trivial to incorporate spatial control in motion generation due to the discrepancy between the relative $\boldsymbol{x}$ and global $\boldsymbol{c}$.
@@ 3.2 Human Motion Diffusion Model (MDM)
Diffusion Process in MDM. Motivated by the success of image diffusion models [22, 50, 70, 9, 53], Motion Diffusion Model (MDM) [55] is proposed to synthesize sequence-level human motions conditioned on texts $\boldsymbol{p}$ via classifier-free guidance [22]. The diffusion process is modeled as a noising Markov process $q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right)=\mathcal{N}\left(\sqrt{\alpha_{t}}\boldsymbol{x}_{t-1},\left(1-\alpha_{t}\right)\boldsymbol{I}\right)$, where $\alpha_{t}\in(0,1)$ are small constant hyper-parameters, thus $\boldsymbol{x}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$ if $\alpha_{t}$ is small enough. Here $\boldsymbol{x}_{t}\in\mathbb{R}^{N\times D}$ is the entire motion sequence at denoising time-step $t$, and there are $T$ time-steps in total. Thus, $\boldsymbol{x}_{0}$ is the clean motion sequence, and $\boldsymbol{x}_{T}$ is almost a random noise to be sampled. The denoising Markov process is defined as $p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{p}\right)=\mathcal{N}\left(\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p}),\left(1-\alpha_{t}\right)\boldsymbol{I}\right),$ where $\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p})$ is the estimated posterior mean for the $t-1$ step from a neural network based on the input $\boldsymbol{x}_{t}$ and $\theta$ is its parameters. Following MDM, we predict the clean motion $\boldsymbol{x}_{0}(\boldsymbol{x}_{t},t,\boldsymbol{p};\theta)$ instead of the noise $\boldsymbol{\epsilon}$ via a transformer [58], and the posterior mean $\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p})$ is
@@ 4 Experiments
Due to the page limit, we put the implementation details, text-to-motion generation and ablation studies in the Appendix B.1, B.2 and B.3.
@@ 5 Conclusion and Limitations
Acknowledgment. This project is funded in part by Shanghai AI Laboratory (P23KS00020, 2022ZD0160201), CUHK Interdisciplinary AI Research Institute, and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. We would like to thank Tianfan Xue for his insightful discussion.
@@ A.1 Pseudo-code of IK guidance
Besides, we also show examples of PyTorch-like codes to implement collision avoidance (1) and making two people face each other (2). By putting such loss function in IK guidance to L-BFGS optimizers, desired orientations or joint distances in motion interactions could be achieved.
def avoid_collision_loss(a_joint, b_joint):
Calculate a loss to avoid collision between two people based on the distance between their torso joints.
Args:
a_joint and b_joint (torch.Tensor): Joint locations of person A in SMPL format, shape [bs, njoints, 3, seqlen].
torso_id = [0, 1, 2, 3, 6, 9, 12, 13, 14, 15]
dim_indices = [0, 2]
a_torso_joints = a_joint[:, torso_id.unsqueeze(1), dim_indices, :]
b_torso_joints = b_joint[:, torso_id.unsqueeze(1), dim_indices, :]
diff = a_torso_joints.unsqueeze(2) - b_torso_joints.unsqueeze(1)
distance = torch.norm(diff, dim=3)
collision_constraint = F.relu(0.4 - distance)
loss = collision_constraint.mean(dim=[1, 2, 3]).mean()
return loss
Listing 1: Avoid Collision Loss
def face_to_face_loss(a_joint, b_joint):
Calculate a loss to make two people face each other based on shoulders and hips orientation and distance.
Args:
a_joint and b_joint (torch.Tensor): Joint locations of person A, shape [bs, njoints, 3, seqlen].
l_hip, r_hip, l_shoulder, r_shoulder = [1, 2, 16, 17]
a_shoulder_vec = a_joint[:, l_shoulder] - a_joint[:, r_shoulder]
a_hip_vec = a_joint[:, l_hip] - a_joint[:, r_hip]
b_shoulder_vec = b_joint[:, l_shoulder] - b_joint[:, r_shoulder]
b_hip_vec = b_joint[:, l_hip] - b_joint[:, r_hip]
a_front_vec = a_joint[:, l_shoulder] - a_joint[:, l_hip]
b_front_vec = b_joint[:, l_shoulder] - b_joint[:, l_hip]
a_shoulder_vec_norm = F.normalize(a_shoulder_vec, dim=2)
a_hip_vec_norm = F.normalize(a_hip_vec, dim=2)
b_shoulder_vec_norm = F.normalize(b_shoulder_vec, dim=2)
b_hip_vec_norm = F.normalize(b_hip_vec, dim=2)
a_front_vec_norm = F.normalize(a_front_vec, dim=2)
b_front_vec_norm = F.normalize(b_front_vec, dim=2)
orientation_loss_shoulder = (a_shoulder_vec_norm + b_shoulder_vec_norm).pow(2).sum(2)
orientation_loss_hip = (a_hip_vec_norm + b_hip_vec_norm).pow(2).sum(2)
front_facing_loss = -torch.sum(a_front_vec_norm * b_front_vec_norm, dim=2)
total_loss = orientation_loss_shoulder.mean() + orientation_loss_hip.mean() + front_facing_loss.mean()
return total_loss
Listing 2: Face-to-Face Loss
Figure 5: Architecture of Motion ControlNet.
@@ A.3 LLM-Planner
A.4 Comparison with a recent Work [65]
We notice that a recent work [65] accepted by ICLR2024 also incorporate the idea of ControlNet [70] and Classifier Guidance [9] in human motion generation. Although it shares similar part of single-person motion control design with our InterControl, we develop our model independently. Here we discuss the detailed difference between our InterControl and their method [65].
  • Main Contribution: We propose the first work to generate multi-person interactions using only single person data for training. It is achieved by our novel insight to treat interactions as contact pairs and transform it to be a controllable single-person motion generation task. Our motivation of generating human interactions is mainly from UniHSI [64], where human-scene interactions could be represented by contacts of human joints and object parts. We extend this idea to human-human interaction by defining human interactions as human joint contact pairs. To leverage such human joint contact pairs for generating human motion interactions, we find that spatially controlling every human joints is important. So we incorporate the idea of ControlNet [70] and Classifier Guidance [9] in motion generation to perform precise spatial joint control. Our implementation of single-person motion generation design is also partially inspired by GMD [27] that could only control root joint. On the contrary, the recent work [65] focus on single-person motion generation without considering any interaction. Its motivation is to extend root control in GMD [27] to any joint control. Besides, we use single-person motion generation model and human joint contact pairs to generate human interactions instead of using few-shot two-people interaction data for training like PriorMDM [51]. Because we do not want the interaction generation ability is learned for fixed number of people. For example, a model treated two-people interaction as a whole data for training could only perform interactions between two people and is unable to generate three-people interactions.
  • LLM-Planner: To convert the interaction descriptions to be human joint contact pairs, we follow UniHSI [64] to leverage LLM-Planners to generate them for us. The recent work [65] focus on single-person motion generation and does not has LLM-Planner.
  • IK guidance: We use L-BFGS [37] for precise alignment of human joint and global condition, which is a standard practice in Inverse Kinematics. It is inspired by the regularization losses in MDM [55] which uses Forward Kinematics. Thus, we believe Inverse Kinematics will be effective for spatial alignment. The recent work [65] uses gradient to align human joints with global conditions, which is a standard practice in classifier guidance [9] in image domain. To facilitate the interactions, we also use two types of joint relations to pull or push the joint distances between interacted humans, while the recent work [65] only consider one joint relation, i.e., be close to spatial condition. We also utilize the guidance on more general form of loss function that describe the human orientations to make people face each other or face the same direction. Yet, The recent work [65] does not consider optimizing orientation.
  • Motion ControlNet: We design extra spatial information $\boldsymbol{c}^{final}$ as spatial condition for motion controlnet, while the recent work [65] directly use original global coordinates of joint positions $\boldsymbol{c}$ as spatial condition. Experimental results show that our spatial information $\boldsymbol{c}^{final}$ performs better than $\boldsymbol{c}$ in controlling all joints.
@@ B.2 Text-to-Motion Generation Results
Table 4: Ablation studies on the HumanML3D [14] dataset.
Item Method FID $\downarrow$ R-precision $\uparrow$
Diversity $\rightarrow$ Foot skating (Top-3)
Traj. err. $\downarrow$ ratio $\downarrow$
Loc. err. $\downarrow$
Avg. err.$\downarrow$
(1) Ours (random joint) 0.178 0.669 9.498 0.0968 0.0403 0.0031 0.0741
(2) w/o ControlNet 0.965 0.621 9.216 0.1624 0.0879 0.0059 0.1013
(4) w/o IK guidance 0.187 0.664 9.598 0.0704 0.8569 0.4553 0.6557
(6) w/ 1-st order grad 0.198 0.668 9.472 0.0987 0.0879 0.0096 0.0877
(7) sparsity = 0.25 0.248 0.671 9.442 0.0801 0.0106 0.0007 0.0546
(8) sparsity = 0.025 0.255 0.663 9.520 0.0705 0.0015 0.0001 0.0067
Table 5: Inference time analysis on a NVIDIA A100 GPU.
Time (s) 39.1 57.3 76.5 80.1
B.3 Ablation Studies
To further investigate the effectiveness of InterControl, we ablate our method in Tab. B.2 and reveal some key information in controlling the motion generation model in the global space. Then we also analyze the computational costs of our method to ensure our control is efficient. We will refer to the variants of InterControl by row numbers in Tab. B.2. All experiments are trained on all joints and evaluated with randomly selected joints to report average performance.
Inference time analysis. In practice, we find that IK guidance in last few denoising steps (e.g., $t\in[0,9]$) is vital for precise joint control, while most denoising steps $t\in[10,999]$ are less important yet take most of computations. IK guidance on $\boldsymbol{x}_{0}$ with only once L-BFGS in $t\in[10,999]$ and $10$ times in $t\in[0,9]$ could leads FID $0.234$ in controlling all joints, yet leads to minimal extra computations. We report its total inference time of $1000$ denoising steps by adding sub-modules step-by-step in Tab. 5. GMD [27] needs $110$s to run two-stage diffusion models, while we only needs $80$s. Gradient-based optimization in the recent work [65] needs $120$s to achieve similar control quality. Thanks to GPU’s parallel computing, generating a batch of $32$ people in InterControl only needs $91$s, enabling efficient group motion generation.
Sparse control signals in temporal. As a key challenge of spatial control is the sparsity, we also report results with sparsely selected frames as control (sparsity = $0.25$ and $0.025$) in Tab. B.2 (row 7 and 8). Our model shows consistent performance of both spatial errors and semantic-level metrics on sparse signals, e.g., FID $0.255$ and avg. err. $0.0467$ with sparsity $0.025$, while GMD [27] achieves FID $0.523$ and avg. err. $0.139$ with the same sparsity.
@@ B.3 Ablation Studies
B.4 More Single-joint Control Results

### v4

@@ Abstract
Text-conditioned motion synthesis has made remarkable progress with the emergence of diffusion models. However, the majority of these motion diffusion models are primarily designed for a single character and overlook multi-human interactions. In our approach, we strive to explore this problem by synthesizing human motion with interactions for a group of characters of any size in a zero-shot manner. The key aspect of our approach is the adaptation of human-wise interactions as pairs of human joints that can be either in contact or separated by a desired distance. In contrast to existing methods that necessitate training motion generation models on multi-human motion datasets with a fixed number of characters, our approach inherently possesses the flexibility to model human interactions involving an arbitrary number of individuals, thereby transcending the limitations imposed by the training data. We introduce a novel controllable motion generation method, InterControl, to encourage the synthesized motions maintaining the desired distance between joint pairs. It consists of a motion controller and an inverse kinematics guidance module that realistically and accurately aligns the joints of synthesized characters to the desired location. Furthermore, we demonstrate that the distance between joint pairs for human-wise interactions can be generated using an off-the-shelf Large Language Model (LLM). Experimental results highlight the capability of our framework to generate interactions with multiple human characters and its potential to work with off-the-shelf physics-based character simulators. Code is available at https://github.com/zhenzhiwang/intercontrol.
@@ 2.3 Controllable Diffusion Models
Diffusion-based generative models have achieved great progress in generating various modalities, such as image [50, 22, 9, 53], video [11, 17, 24] and audio [32]. Conditions and controlling ability in diffusion models are also well studied: (1) Inpainting-based methods [8, 7] predict part of the data with the observed parts as condition and rely on diffusion model to generate consistent output, which is used in PriorMDM [51]. (2) Classifier-guidance [9] trains a separate classifier and exploits the gradient of classifier to guide the diffusion process. Our InterControl inherits the spirit of classifier-guidance, yet our guidance is provided by Inverse Kinematics (IK) and no classifier is needed. (3) Classifier-free guidance [22] trains a conditional and an unconditional diffusion model simultaneously and trade-off its quality and diversity by setting weights. (4) ControlNet [70] introduces a trainable copy of pretrained diffusion model to process the condition and freezes the original model to avoid degeneration of generation ability. It enables diverse types of dense control signals for various purpose with minimal finetuning effort. Our InterControl also incorporate the idea of ControlNet [70] to finetune the pretrained MDM [55] to process spatial control signals and improve the quality of generated motions after joint control.
@@ 3.1 Formulation of Interaction Generation
Inspired by human-scene interaction [64], we define human interactions as joint contact pairs $\mathcal{C}=\left\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots\right\}$, where $\mathcal{S}_{i}$ is the $i^{th}$ contact step. Taking two-person interaction as an example, each step $\mathcal{S}$ has several contact pairs $\mathcal{S}=\left\{\left\{j^{1}_{1},j^{2}_{1},t^{s}_{1},t^{e}_{1},c_{1},d_{1}\right\},\left\{j^{1}_{2},j^{2}_{2},t^{s}_{2},t^{e}_{2},c_{2},d_{2}\right\},\ldots\right\}$, where $j^{1}_{k}$ is the joint of person 1, $j^{2}_{k}$ is the joint of person 2, $t^{s}_{k}$ and $t^{e}_{k}$ means the start and end frame of the interaction, $c_{k}$ means contact type from {contact, avoid} to pull or push the joint pairs, $d_{k}$ is the desired distance in the interaction. By converting the contact pairs $\mathcal{S}$ to the mask $\boldsymbol{m}$ and distance $d$, and taking others’ joint positions as condition, we could guide the multi-person motion generation process to interact between joints in the form of spatial distance. In this way, interaction generation is transformed to be controllable single-person motion generation taking a text prompt $\boldsymbol{p}$ and a spatial control signal $\boldsymbol{c}\in\mathbb{R}^{N\times J\times 3}$ as input. Its goal is to predict motion sequence $\boldsymbol{x}\in\mathbb{R}^{N\times D}$ whose joints in the global space is aligned with spatial control $\boldsymbol{c}$, where $N$ is number of frames, $J$ is number of joints (e.g., 24 in SMPL [38]), and $D$ is the dimension of relative joint representations (e.g., 263 in HumanML3D [14]). Incorporating spatial control in motion generation presents challenges due to the discrepancy between relative motion representation $\boldsymbol{x}$ and global $\boldsymbol{c}$.
@@ 3.2 Human Motion Diffusion Model (MDM)
Diffusion Process in MDM. Motivated by the success of image diffusion models [22, 50, 70, 9, 53], Motion Diffusion Model (MDM) [55] is proposed to synthesize sequence-level human motions conditioned on texts $\boldsymbol{p}$ via classifier-free guidance [22]. The diffusion process is modeled as a noising Markov process $q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right)=\mathcal{N}\left(\sqrt{\alpha_{t}}\boldsymbol{x}_{t-1},\left(1-\alpha_{t}\right)\boldsymbol{I}\right)$, where $\alpha_{t}\in(0,1)$ are small constant hyper-parameters, thus $\boldsymbol{x}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$ if $\alpha_{t}$ is small enough. Here $\boldsymbol{x}_{t}\in\mathbb{R}^{N\times D}$ is the entire motion sequence at denoising time-step $t$, and there are $T$ time-steps in total. Thus, $\boldsymbol{x}_{0}$ is the clean motion sequence, and $\boldsymbol{x}_{T}$ is a random noise to be sampled. The denoising Markov process is defined as $p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{p}\right)=\mathcal{N}\left(\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p}),\left(1-\alpha_{t}\right)\boldsymbol{I}\right),$ where $\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p})$ is the estimated posterior mean for the $t-1$ step from a neural network based on the input $\boldsymbol{x}_{t}$ and $\theta$ is its parameters. Following MDM, we predict the clean motion $\boldsymbol{x}_{0}(\boldsymbol{x}_{t},t,\boldsymbol{p};\theta)$ instead of the noise $\boldsymbol{\epsilon}$ via a transformer [58], and the posterior mean $\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{p})$ is
@@ 4 Experiments
Due to the page limit, we put the implementation details and text-to-motion generation in the Appendix B.1 and B.2.
@@ 4.2 Zero-Shot Multi-Person Interaction Generation
Table 3: Ablation studies on the HumanML3D [14] dataset.
Item Method FID $\downarrow$ R-precision $\uparrow$
Diversity $\rightarrow$ Foot skating (Top-3)
Traj. err. $\downarrow$ ratio $\downarrow$
Loc. err. $\downarrow$
Avg. err.$\downarrow$
(1) Ours (random joint) 0.178 0.669 9.498 0.0968 0.0403 0.0031 0.0741
(2) w/o ControlNet 0.965 0.621 9.216 0.1624 0.0879 0.0059 0.1013
(4) w/o IK guidance 0.187 0.664 9.598 0.0704 0.8569 0.4553 0.6557
(6) w/ 1-st order grad 0.198 0.668 9.472 0.0987 0.0879 0.0096 0.0877
(7) sparsity = 0.25 0.248 0.671 9.442 0.0801 0.0106 0.0007 0.0546
(8) sparsity = 0.025 0.255 0.663 9.520 0.0705 0.0015 0.0001 0.0067
Table 4: Inference time analysis on a NVIDIA A100 GPU.
Time (s) 39.1 57.3 76.5 80.1
4.3 Ablation Studies
To further investigate the effectiveness of InterControl, we ablate our method in Tab. 4.2 and reveal some key information in controlling the motion generation model in the global space. Then we also analyze the computational costs of our method to ensure our control is efficient. We will refer to the variants of InterControl by row numbers in Tab. 4.2. All experiments are trained on all joints and evaluated with randomly selected joints to report average performance.
Inference time analysis. In practice, we find that IK guidance in last few denoising steps (e.g., $t\in[0,9]$) is vital for precise joint control, while most denoising steps $t\in[10,999]$ are less important yet take most of computations. IK guidance on $\boldsymbol{x}_{0}$ with only once L-BFGS in $t\in[10,999]$ and $10$ times in $t\in[0,9]$ could leads FID $0.234$ in controlling all joints, yet leads to minimal extra computations. We report its total inference time of $1000$ denoising steps by adding sub-modules step-by-step in Tab. 4. GMD [27] needs $110$s to run two-stage diffusion models, while we only needs $80$s. Gradient-based optimization in the recent work [65] needs $120$s to achieve similar control quality. Leveraging GPU parallel computing capabilities, InterControl can efficiently generate motions for a batch of $32$ people in $91$ seconds, enabling efficient group motion generation.
Sparse control signals in temporal. As a key challenge of spatial control is the sparsity, we also report results with sparsely selected frames as control (sparsity = $0.25$ and $0.025$) in Tab. 4.2 (row 7 and 8). Our model demonstrates consistent performance in both spatial error and semantic-level metrics when using sparse signals, e.g., FID $0.255$ and avg. err. $0.0467$ with sparsity $0.025$, while GMD [27] achieves FID $0.523$ and avg. err. $0.139$ with the same sparsity.
@@ 5 Conclusion and Limitations
Acknowledgment. This project is funded in part by Shanghai Artificial Intelligence Laboratory, CUHK Interdisciplinary AI Research Institute, and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. We would like to thank Tianfan Xue for his insightful discussion.
@@ B.2 Text-to-Motion Generation Results
B.3 More Single-joint Control Results